Find the exact value of sin(a - b) if sin a = 15 12 17 and sin B = and a and Bare between 0 and 1/2. 13
Answer #1

\text{Answer: }~~\sin(\alpha -\beta)= -{21\over 221}

Sin (a-ß) = sin(a) 6 (P) - Go (d) Sin (A) Given, Sina= 15 17 & Sinß = 12 13 where a 4 ß are between of TT 2 we know, sino + C